An interesting (but oversimplified) model of an atom pictures an electron “in orbit” around a proton. Suppose this electron is moving in a circular orbit of radius 0.10 nm (1.0 x 10-10) and the force that makes this circular motion possible is the electric force exerted by the proton on the electron. Find the speed of the electron.
1
Expert's answer
2020-05-26T12:54:47-0400

Let us calculate the electric force exerted by the proton on the electron:

"F = \dfrac{k_eq_pq_e}{r^2}" . Module of charges of proton and electron is "q = 1.6\cdot10^{-19}\,\mathrm{C}." Therefore, the force is

"F = \dfrac{9\cdot10^9\cdot (1.6\cdot10^{-19})^2}{(1.0\cdot10^{-10})^2} = 2.3\cdot10^{-8}\,\mathrm{N}." The acceleration can be calculated as

"a = \dfrac{F}{m_e}" , but also the acceleration is "a = \dfrac{v^2}{r}." Therefore, velocity is

"v = \sqrt{\dfrac{Fr}{m_e}} = \sqrt{\dfrac{2.3\cdot10^{-8}\,\mathrm{N}\cdot1.0\cdot10^{-10}\,\mathrm{m}}{9.1\cdot10^{-31}\,\mathrm{kg}}} = 1.6\cdot10^3\,\mathrm{km/s}."